OpenStudy (anonymous):

Solve for x. 2x^2-4x+1=0

OpenStudy (lgbasallote):

use quadratic formula \[\large x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

Is it an equation or simply an expression??? meaning: \[2x^2 - 4x + 1 \qquad or \qquad 2x^2 - 4x + 1 = 0\quad ?\]

OpenStudy (anonymous):

I'm stuck at x = (8+2sqrt2)/4, (8-2sqrt2)/4... is this the simplest form?

OpenStudy (anonymous):

fololl lgbasallote formula , a=2 , b= -4 , c=1

OpenStudy (anonymous):

Tell me what is a, b and c here by comparing it with standard equation: \[ax^2 + bx + c = 0\]

OpenStudy (anonymous):

can u solve?

OpenStudy (lgbasallote):

i think he just did lol

OpenStudy (lgbasallote):

@derptastic878 factor out

OpenStudy (lgbasallote):

im not sure if you're right but i'll assume you are \[x = \frac{8 + 2\sqrt 2}{4}\] \[x = \frac{2(4 + \sqrt 2)}{4}\] \[x = \frac{4 + \sqrt 2}{2}\] do the same for the other one

OpenStudy (anonymous):

Thanks!

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (anonymous):

Can x = (8-2sqrt2)/4 be factored? The 2 is negative.

OpenStudy (lgbasallote):

yes it can :)

OpenStudy (anonymous):

2(4-sqrt2) ?

OpenStudy (lgbasallote):

yes that's the numerator

OpenStudy (anonymous):

2(4-sqrt2)/2 :D?